# Federated Foundation Model for Cardiac CT Imaging

### v1

@@ 4 Methods
This study’s data comprises patients who underwent a minimally invasive procedure for replacing their aortic valve with a Transcatheter Aortic Valve Implantation (TAVI) prosthesis. Consistent with clinical guidelines, every patient undergoes a contrast-enhanced CT scan triggered by an electrocardiogram, conducted in either only the systolic or both the systolic and diastolic phases of the heart cycle. For this study we included all available contrast enhanced CT scans not dependent whether they only had the systolic or diastolic phase available. Collective information about the demographics of the included population and CT imaging parameters is presented in Fig. 7.
Figure 7: Demographics of patients and data properties across clients. Some data was not available at all clients. Three manufacturers with in total eleven different models were included in the federated training. The acquisition protocols in terms of exposure, exposure time, X-ray tube current, and contrast bolus volume vary across clients. Manufacture acronyms are P: Philips, S: Siemens, T: Toshiba.
The data acquisition occurred at each participating site from 2015 to 2021. Each site’s institutional review board approved the retrospective analysis of CT scans from patients who received a TAVI prosthesis during this time. However, challenges in exporting data from the PACS varied by location, preventing the complete dataset from being utilized for model training or testing at some sites. These challenges primarily involved limitations in automatically exporting large volumes of data from the internal PACS systems. Our study highlighted deficiencies in data export protocols at some institutions, which we hope will trigger investments into better data pipelines. Future studies leveraging this infrastructure can benefit from the insights we have gained.
@@
Where not already present, the annotations (3D points for hinge points, origins of coronary arteries, and membranous septum, and segmentations for calcification) were obtained with the medical Medical Interaction Toolkit (MITK) [^3] https://www.mitk.org/wiki/The_Medical_Imaging_Interaction_Toolkit_(MITK). Annotation protocols were provided in text and video form, which was reported to be very beneficial for uniform label generation.
@@
For the three subtasks we used the popular 3D UNet with residual connections (3D-ResUNet) with 32 base filters [8, 13, 15]. The learning rate was set to $lr=0.01$ and optimized with the AdamW optimizer [27]. As loss function during training we used a combination of cross entropy and DICE score loss with deep supervision [48]. When applying deep supervision also for the intermediate outputs of the skip connections the loss function is applied to a downsampled version of the target, which has been shown to improve segmentation performance [15].
For the final model that combines the knowledge from the three subnetworks we use the new Swin UNet Transformer (Swin UNETR) [12]. We use a feature size of 24 with a patch size of $\mathbb{R}^{96\times 96\times 96}$. The learning rate was set to $lr=10^{-4}$ and optimized with the AdamW optimizer [27]. We equipped the transformer with three heads, one for each task, to train all tasks concurrently.
@@
In federated learning multiple data holding clients train a model locally on their data shards and report the trained model weights back to a central server where averaging is performed [30, 19]. After successful averaging another round of training is initiated until the model converges. Each round is termed a federated round. This allows data privacy compliant model training as no patient data ever leaves the individual hospitals boundaries. The most widespread architecture is a hub-and-spoke system were all clients train in parallel instead of an e.g. sequential training [18, 36].
Our federation spans eight cardiology and radiology department in university hospitals in Germany (c.f. Fig. 1). Connection could be established only from within the individual clinics to a server that resided behind a firewall at Heidelberg University. Each model was trained for 20 federated rounds of averaging with 10 local epochs in each round. We chose to perform model weight’s aggregation using a popular variant of the federated averaging algorithm [23]. Every communication in our federation was based on transport layer security, additional authentication with username and password, and server-side IP address white listing. These measures help mitigate some of the privacy and security concerns still inherent to FL.
Our work covers the whole process of extracting real world data from clinical information systems and subsequent homogenization of data formats across the different sites and label types. The federated learning software stack was installed at each location that is intended to be used beyond this study for future research. We created a custom fork of the renowned Kaapana platform [44]. It allows for a flexible deployment of containerized applications in combination to a picture archiving and communication system (PACS). To extract the cohorts needed per location we use a custom developed filtering tool described here [53]. Each data type is stored in a custom structured report template such that they can be linked to the corresponding series [53]. Segmentation objects can also be stored and linked to the referenced image series within the PACS. Fedbiomed is used as FL library as they provide very sophisticated security measures [47]. All communication is encrypted with transport layer security (TLS) encryption, where the key is distributed to the clients prior to training. Further, each client must authenticate with custom credentials (username and password). And last, IP white listing is performed such that only predefined IP addresses can initiate a connection. The connection is unidirectional. It must be initiated from within the clinic network, the clients then poll for updates such that no action can be triggered from the server without the client noticing.
@@
All data from the eight sites used in this study are not made publicly available due to restrictions imposed by the participating sites. The data was also not publicly available during conducting of this study. As by privacy-by-design definition of federated learning they were instead used locally during training and validation of the trained models. The data to reproduce the plots as well as the corresponding scripts are made publicly available under: https://github.com/Cardio-AI/fed-foundation-model-cardiac-ct. The ImageCAS dataset is available under: https://github.com/XiaoweiXu/ImageCAS-A-Large-Scale-Dataset-and-Benchmark-for-Coronary-Artery-Segmentation-based-on-CT. The corresponding labels for quantifying the inter-observer variability are available at: https://github.com/Cardio-AI/fed-foundation-model-cardiac-ct. The pointsets can be opened with the Medical Interaction Toolkit (MITK) available under: https://www.mitk.org/wiki/The_Medical_Imaging_Interaction_Toolkit_(MITK).
@@
Following the FAIR criteria (findability, accessibility, interoperability, and reusability) in scientific research all code used in this study is made publicly available. We used a custom fork of Kaapana [44] from https://github.com/kaapana/kaapana which is available under https://github.com/Cardio-AI/kaapana for orchestration of docker containers at each location. The federeated learning library fedbiomed is available under https://github.com/fedbiomed/fedbiomed our custom fork with more security features enabled is avilable under https://github.com/Cardio-AI/fedbiomed. For creation of labels we use MITK https://www.mitk.org/wiki/The_Medical_Imaging_Interaction_Toolkit_(MITK). The nnUNet pipeline used for training the per-task models is available under https://github.com/MIC-DKFZ/nnUNet. Our preprocessing, training, and validation scripts are made available unde https://github.com/Cardio-AI/fed-foundation-model-cardiac-ct. The pipelines were developed using PyTorch [34], MONAI [7], TorchIO [37], and SimpleITK [4].
@@ Acknowledgements
Malte Tölle developed the presented method, conducted the experiments and evaluations, set up the federated learning software stack, and wrote the manuscript. Sandy Engelhardt organised the consortium, and significantly helped to shape the methods and the manuscript. Sebastian Eble, Moritz Bednorz, and Halvar Kelm contributed code for model training and software setup. Philipp Garthe, Lars Kaderali, Nina Krüger, Andreas Leha, Simon Martin, Clemens Scheree, Jan Moritz Seliger, and Stefan Simm were the direct contact persons at each participating location, set up the local infrastructures, and exported, curated, and uploaded the required data. Florian Andre, Peter Bannas, Norbert Frey, Stefan Groß, Alexander Meyer, Eike Nagel, Stefan Orwat, Tim Friede, Tim Seidler, and Sandy Engelhardt developed the idea for the study, provided guidance, and helped with revising the final manuscript.
@@
Norbert Frey reports speaker honoraria, presentations or advisory board consultations from AstraZeneca, Bayer AG, Boehringer Ingelheim, Novartis, Pfizer, Daiichi Sankyo Deutschland. Tim Seidler reports research, educational, or travel grants and honoraria for lectures or advisory board consultations from Abbott Vascular, AstraZeneca, BoehringerIngelheim, Bristol Myers Squibb, Corvia, Cytokinetics, Edwards Life Sciences, Medtronic, Myocardia, Novartis, Pfizer, Teleflex. Alexander Meyer reports consulting or lecturing fees from Medtronic, Bayer, Pfizer. Clemens Scherer reports speaker honorarium from AstraZeneca. Sandy Engelhardt reports speaker honorarium from Boehringer Ingelheim. None are related to the content of the manuscript. The other authors declare no conflicts of interest.

### v2

@@ Methods
Figure 6: Data distribution across locations. In total $8,104$ CT scans are available across all eight locations. For each label the distribution is differently skewed. While the most uniform distribution is present for the hinge point training, for membranous septum and calcification the distribution is skewed. Still, the federated model that is trained over these skewed distributions exhibits better performance than the one trained on a single client. HPs & COs: Hinge Points and Coronary Ostia Points, MS: Membranous Septum.
@@
All procedures performed in studies involving human participants were in accordance with the ethical standards of the institutional and/or national research committee and with the 1964 Helsinki declaration and its later amendments or comparable ethical standards. This article does not contain any studies with animals performed by any of the authors. Ethical approval was waived by the local Ethics Committees of Heidelberg (S-475/2021), Göttingen (11/6/21), Hamburg (2021-200262-BO-bet), Munich (21-0497), Münster (2021-487-b-S), Greifswald (BB 091/24), and Frankfurt (2021-366_1) in view of the retrospective nature of the study and all the procedures being performed were part of the routine care. In Berlin, a multi-centric study must not explicitly be confirmed when another institutional ethics board waived approval.
This study’s data comprises patients who underwent a minimally invasive procedure for replacing their aortic valve with a Transcatheter Aortic Valve Implantation (TAVI) prosthesis. Consistent with clinical guidelines, every patient undergoes a contrast-enhanced CT scan triggered by an electrocardiogram, conducted in either only the systolic or both the systolic and diastolic phases of the heart cycle. For this study we included all available contrast enhanced CT scans not dependent whether they only had the systolic or diastolic phase available. Collective information about the demographics of the included population and CT imaging parameters is presented in Supplementary Figure 1.
The data acquisition was performed at each participating site from 2015 to 2021. Each site’s institutional review board approved the retrospective analysis of CT scans from patients who received a TAVI prosthesis during this time. However, challenges in exporting data from the PACS varied by location, preventing the complete dataset from being utilized for model training or testing at some sites. These challenges primarily involved limitations in automatically exporting large volumes of data from the internal PACS systems. Our study highlighted deficiencies in data export protocols at some locations, which we hope will trigger investments into better data pipelines. Future studies leveraging this infrastructure can benefit from the insights we have gained.
Training is performed on the data quantities across locations as shown in Figure 6. In total, we have 8104 CT images (all locations), 251 hinge points and coronary ostia (HPs & COs) (7 locations), 116 membranous septum (MS) (3 locations), and 118 calcification labels (3 locations). None of the displayed data distributions are uniform. Location (L) 3 has the highest number of CT images (3001), while L5 has only 20. Seven locations have HPs & COs labeled, with a maximum of 78 cases and a minimum of 20. Three locations contain labels for membranous septum (73/20) and calcification (78/20). The sample heterogeneity is notably large, especially in comparison to previous studies in the field [8, 10, 12]. Additionally, no two locations have similar distributions of images or label types.
@@
Where not already present, the annotations (3D points for hinge points, origins of coronary arteries, and membranous septum, and segmentations for calcification) were obtained with the medical Medical Interaction Toolkit (MITK) [71]. Annotation protocols were provided in text and video form, which was reported to be very beneficial for uniform label generation.
Semi-supervised Federated Knowledge Distillation
Figure 7: Overview of two-stage federated learning process with knowledge distillation. a) In stage 1 one network per label type is trained with FL across all locations in possession of that particular label type (A). These trained models are used to create pseudo labels on the unlabeled datasets residing in each hospital (B). b) In stage 2, a transformer based architecture is trained on the generated pseudo labels (A). The model gains one prediction head per label type. Finally, the model’s prediction heads are finetuned on the ground truth labels while leaving the backbone’s weights fixed (B). Each label type is visualized with a distinct color (blue, yellow, orange), unlabeled data is shown in gray. If pseudo-labels are generated for the previously unlabeled data with the models from stage 1, it is marked with the three label colors.
Our proposed semi-supervised, two-stage federated learning approach enables effective training on large datasets by leveraging unlabeled data. In the first stage, we train a convolutional model on labeled subsets of the data. Since some locations only have partial labels, we train a separate model for each label type (i.e., hinge points and coronary arteries, membranous septum, and calcification). These three specialized models generate pseudo-labels on the unlabeled data at each site.
In the second stage, we use these pseudo-labels to train a transformer model with a unified structure that includes a prediction head for each label type. This setup combines knowledge across the three tasks into one model through a semi-supervised, federated knowledge distillation (KD) process, condensing the knowledge of three models into a single, comprehensive model. Figure 7 visualizes the two-stage training procedure.
Finally, we finetune each prediction head on the labeled data at each client, while keeping the weights of the shared feature extractor fixed. This design allows the model to learn task-agnostic features in the backbone, meaning features for different segmentation and localization tasks, like identifying the aortic root or the aortic valve hinge points, are captured in a unified manner.
@@
For the three subtasks we used the popular 3D UNet with residual connections (3D-ResUNet) with 32 base filters [18, 72, 73]. The learning rate was set to $lr=0.01$ and optimized with the AdamW optimizer [74]. As loss function during training we used a combination of cross entropy and DICE score loss with deep supervision [75]. When applying deep supervision also for the intermediate outputs of the skip connections the loss function is applied to a downsampled version of the target, which has been shown to improve segmentation performance [18]. For guidance we feed the segmented heart obtained from the TotalSegmentator tool [70] as a condition such that the models can learn the anatomical relations between heart and the corresponding structures.
For the final model that combines the knowledge from the three subnetworks we use the SWIN UNet Transformer (SWIN-UNETR) [43]. We use a feature size of 24 with a patch size of $\mathbb{R}^{96\times 96\times 96}$. The learning rate was set to $lr=10^{-4}$ and optimized with the AdamW optimizer [74]. We equipped the transformer with three heads, one for each task, to train all tasks concurrently. We again add the heart segmentation as input for anatomical guidance. For comparison we also train a conventional vision transformer based segmentation model [42]. We used a hidden dimension of 768 and a patch size of $\mathbb{R}^{16\times 16\times 16}$. Optimizer setting and inputs are similar to the SWIN-UNETR. The transformer based architectures vary in their attention mechanism. While the ViT employs conventional self-attention, the SWIN-UNETR uses shifted window self-attention.
@@
In federated learning multiple data holding locations train a model locally on their data shards and report the trained model weights back to a central server where averaging is performed [76, 77]. After successful averaging another round of training is initiated until the model converges. Each round is termed a federated round. This allows data privacy compliant model training as no patient data ever leaves the individual hospitals boundaries. The most widespread architecture is a hub-and-spoke system were all locations train in parallel instead of an e.g. sequential training [5, 45].
Our federation spans eight cardiology and radiology department in university hospitals in Germany (c.f. Figure 1). Connection could be established only from within the individual clinics to a server that resided behind a firewall at Heidelberg University. Each model was trained for 20 federated rounds of averaging with 10 local epochs in each round. We chose to perform model weight’s aggregation using a popular variant of the federated averaging algorithm [78]. Every communication in our federation was based on transport layer security, additional authentication with username and password, and server-side IP address white listing. These measures help mitigate some of the privacy and security concerns still inherent to FL.
Our work covers the whole process of extracting real world data from clinical information systems and subsequent homogenization of data formats across the different sites and label types. The federated learning software stack was installed at each location that is intended to be used beyond this study for future research. We created a custom fork of the renowned Kaapana platform [57]. It allows for a flexible deployment of containerized applications in combination to a picture archiving and communication system (PACS). To extract the cohorts needed per location we use a custom developed filtering tool [58]. Each data type is stored in a custom structured report template such that they can be linked to the corresponding series. Segmentation objects can also be stored and linked to the referenced image series within the PACS. Fedbiomed is used as FL library as they provide very sophisticated security measures [56]. All communication is encrypted with transport layer security (TLS) encryption, where the key is distributed to the locations prior to training. Further, each client must authenticate with custom credentials (username and password). And last, IP white listing is performed such that only predefined IP addresses can initiate a connection. The connection is unidirectional. It must be initiated from within the clinic network, the locations then poll for updates such that no action can be triggered from the server without the client noticing.
@@
All data from the eight sites used in this study are not made publicly available due to restrictions imposed by the participating sites. The data was also not publicly available during conducting of this study. As by privacy-by-design definition of federated learning they were instead used locally during training and validation of the trained models. The data to reproduce the plots as well as the corresponding scripts are made publicly available under: https://github.com/Cardio-AI/FedKD-for-Cardiac-CT. The ImageCAS dataset is available under: https://github.com/XiaoweiXu/ImageCAS-A-Large-Scale-Dataset-and-Benchmark-for-Coronary-Artery-Segmentation-based-on-CT. The corresponding labels for quantifying the inter-observer variability are available at: https://github.com/Cardio-AI/FedKD-for-Cardiac-CT. The pointsets can be opened with the Medical Interaction Toolkit (MITK) available under: https://www.mitk.org/wiki/The_Medical_Imaging_Interaction_Toolkit_(MITK).
@@
Following the FAIR criteria (findability, accessibility, interoperability, and reusability) in scientific research all code used in this study is made publicly available. We used a custom fork of Kaapana [57] from https://github.com/kaapana/kaapana which is available under https://github.com/Cardio-AI/kaapana for orchestration of docker containers at each location. The federeated learning library fedbiomed is available under https://github.com/fedbiomed/fedbiomed our custom fork with more security features enabled is avilable under https://github.com/Cardio-AI/fedbiomed. For creation of labels we use MITK https://www.mitk.org/wiki/The_Medical_Imaging_Interaction_Toolkit_(MITK). The nnUNet pipeline used for training the per-task models is available under https://github.com/MIC-DKFZ/nnUNet. Our preprocessing, training, and validation scripts are made available under https://github.com/Cardio-AI/FedKD-for-Cardiac-CT. The pipelines were developed using PyTorch [79], MONAI [80], TorchIO [81], and SimpleITK [82].
@@ Acknowledgements
MT developed the presented method, conducted the experiments and evaluations, set up the federated learning software stack, and wrote the manuscript. SE organised the consortium, and significantly helped to shape the methods and the manuscript. SEb, MB, and HK contributed code for model training and software setup. PG, LK, NK, AL, SM, CS, JMS, and SS were the direct contact persons at each participating location, set up the local infrastructures, and exported, curated, and uploaded the required data. FA, PB, GD, NF, SG, AH, AM, EN, SO, MS, TF, TS, and SE developed the idea for the study, provided guidance, and helped with revising the final manuscript.
@@
NF reports speaker honoraria, presentations or advisory board consultations from AstraZeneca, Bayer AG, Boehringer Ingelheim, Novartis, Pfizer, Daiichi Sankyo Deutschland. TS reports research, educational, or travel grants and honoraria for lectures or advisory board consultations from Abbott Vascular, AstraZeneca, BoehringerIngelheim, Bristol Myers Squibb, Corvia, Cytokinetics, Edwards Life Sciences, Medtronic, Myocardia, Novartis, Pfizer, Teleflex. AM reports consulting or lecturing fees from Medtronic, Bayer, Pfizer. CS reports speaker honorarium from AstraZeneca. SE reports speaker honorarium from Boehringer Ingelheim. None are related to the content of the manuscript. The authors declare no conflicts of interest.
Supplementary Information
Supplementary Figure 1: Demographics of patients and data properties across locations. Some data was not available at all locations. Three manufacturers with in total eleven different models were included in the federated training. The acquisition protocols in terms of exposure, exposure time, X-ray tube current, and contrast bolus volume vary across locations. Manufacture acronyms are P: Philips, S: Siemens, T: Toshiba.
Supplementary Table 1: Results of local, federated, and knowledge distilled models per location for the task of detecting hinge points and coronary ostia (HPs & COs). ed and KD are trained on L1,2,3,4,6. The local models often overfit to the training data and even underperform on their respective testset. The federated and especially knowledge distilled models show better generalization. All values are reported in mm with mean and standard deviation.
Train L1 L2 L3 L4 L5 L6 L7
UNet L1 $3.58\pm 5.17$ $3.74\pm 2.03$ $3.67\pm 1.68$ $4.35\pm 3.94$ $3.23\pm 1.83$ $2.85\pm 1.62$ $4.84\pm 3.46$
L2 $17.15\pm 21.42$ $11.65\pm 12.59$ $11.24\pm 11.08$ $12.31\pm 9.99$ $12.63\pm 12.29$ $11.71\pm 11.43$ $13.56\pm 11.62$
L3 $5.98\pm 7.54$ $4.79\pm 3.2$ $5.19\pm 8.77$ $5.01\pm 3.52$ $4.88\pm 2.17$ $3.77\pm 2.03$ $5.35\pm 3.72$
L4 $4.81\pm 5.35$ $4.9\pm 3.19$ $8.0\pm 10.94$ $4.53\pm 4.05$ $4.14\pm 2.3$ $3.81\pm 5.95$ $5.44\pm 3.84$
L6 $4.58\pm 3.75$ $4.83\pm 3.74$ $4.97\pm 3.22$ $4.53\pm 3.50$ $3.29\pm 4.32$ $3.12\pm 3.96$ $4.73\pm 3.88$
Fed $3.92\pm 5.6$ $3.53\pm 2.36$ $3.58\pm 1.79$ $4.12\pm 4.0$ $3.41\pm 4.18$ $2.71\pm 1.28$ $3.86\pm 3.09$
KD $3.84\pm 5.3$ $3.77\pm 1.93$ $3.26\pm 1.78$ $4.43\pm 4.03$ $3.37\pm 1.83$ $2.61\pm 1.24$ $4.59\pm 3.77$
ViT L1 $4.93\pm 5.23$ $4.7\pm 2.48$ $3.82\pm 2.11$ $6.1\pm 3.71$ $5.1\pm 9.43$ $3.76\pm 2.16$ $5.75\pm 3.61$
L2 $18.92\pm 21.33$ $15.45\pm 17.23$ $14.63\pm 13.21$ $14.99\pm 12.15$ $13.72\pm 12.30$ $15.78\pm 17.48$ $13.78\pm 14.63$
L3 $15.62\pm 26.09$ $8.65\pm 14.66$ $7.6\pm 14.63$ $13.58\pm 21.33$ $12.68\pm 20.17$ $9.49\pm 14.03$ $16.03\pm 24.19$
L4 $26.87\pm 7.05$ $25.24\pm 6.42$ $24.65\pm 5.02$ $24.75\pm 10.91$ $26.06\pm 6.18$ $23.39\pm 6.5$ $26.33\pm 5.73$
L6 $5.37\pm 34.83$ $5.11\pm 3.98$ $5.12\pm 4.87$ $5.62\pm 3.28$ $4.05\pm 5.45$ $6.36\pm 4.05$ $5.82\pm 3.96$
Fed $4.68\pm 5.26$ $4.47\pm 2.62$ $3.75\pm 2.18$ $5.4\pm 3.81$ $3.87\pm 1.97$ $3.09\pm 1.28$ $5.11\pm 3.53$
KD $4.68\pm 5.25$ $4.4\pm 2.57$ $3.69\pm 2.14$ $5.37\pm 3.81$ $3.9\pm 1.94$ $3.07\pm 1.42$ $5.01\pm 3.43$
SWIN L1 $3.61\pm 6.17$ $3.62\pm 2.22$ $3.76\pm 1.72$ $5.8\pm 7.03$ $2.76\pm 1.47$ $2.79\pm 1.7$ $4.29\pm 3.28$
L2 $15.93\pm 28.61$ $3.9\pm 2.01$ $7.02\pm 8.66$ $17.9\pm 25.69$ $15.55\pm 20.56$ $17.13\pm 12.91$ $11.47\pm 19.15$
L3 $15.86\pm 13.07$ $9.99\pm 10.17$ $3.65\pm 5.66$ $14.09\pm 9.64$ $8.2\pm 9.96$ $9.21\pm 11.0$ $10.49\pm 11.28$
L4 $4.66\pm 7.14$ $3.54\pm 2.27$ $3.48\pm 1.81$ $4.84\pm 6.42$ $3.58\pm 1.83$ $1.97\pm 1.02$ $4.17\pm 3.25$
L6 $5.12\pm 6.23$ $3.58\pm 4.83$ $3.21\pm 2.91$ $4.58\pm 3.27$ $2.86\pm 2.75$ $3.42\pm 2.53$ $3.92\pm 3.33$
Fed $4.73\pm 7.16$ $3.47\pm 1.96$ $3.93\pm 1.85$ $5.65\pm 6.63$ $3.02\pm 1.68$ $2.85\pm 1.49$ $4.34\pm 3.17$
KD $3.49\pm 5.36$ $3.26\pm 1.93$ $3.18\pm 1.78$ $4.17\pm 3.94$ $2.94\pm 1.83$ $2.39\pm 1.11$ $4.11\pm 3.35$
Supplementary Table 2: Results of local, federated, and knowledge distilled models per location for the task of detecting the membranous septum (MS). Fed and KD are trained on L1 and L3. The local models sometimes overfit to the training data and even underperform on their respective testset. The federated and especially knowledge distilled models show better generalization. All values are reported in mm with mean and standard deviation.
Train L1 L3 L7
UNet L1 $3.45\pm 2.63$ $5.10\pm 0.60$ $5.01\pm 2.33$
L3 $4.68\pm 2.73$ $3.66\pm 1.06$ $4.36\pm 1.88$
Fed $4.64\pm 2.33$ $3.72\pm 1.34$ $4.37\pm 2.41$
KD $3.26\pm 2.34$ $3.25\pm 1.32$ $3.40\pm 1.56$
ViT L1 $3.55\pm 2.55$ $3.29\pm 1.53$ $4.26\pm 2.65$
L3 $54.28\pm 36.64$ $24.52\pm 18.99$ $53.98\pm 34.95$
Fed $3.69\pm 2.54$ $4.49\pm 1.91$ $5.39\pm 2.64$
KD $3.34\pm 2.39$ $2.97\pm 1.50$ $3.60\pm 1.56$
SWIN L1 $4.44\pm 3.55$ $4.75\pm 1.98$ $4.92\pm 1.63$
L3 $3.94\pm 2.33$ $3.04\pm 0.91$ $4.60\pm 2.31$
Fed $3.17\pm 2.43$ $3.30\pm 1.60$ $3.43\pm 1.44$
KD $3.29\pm 2.44$ $2.72\pm 0.96$ $3.29\pm 1.45$
Supplementary Table 3: Results of local, federated, and knowledge distilled models per location for the task of segmenting the calcification. Fed and KD are trained on L1 and L2. The DICE scores are reported with mean and standard deviation.
Train L1 L2 L6
UNet L1 $0.593\pm 0.233$ $0.539\pm 0.134$ $0.583\pm 0.412$
L2 $0.391\pm 0.170$ $0.401\pm 0.207$ $0.272\pm 0.274$
Fed $0.486\pm 0.193$ $0.515\pm 0.246$ $0.391\pm 0.212$
KD $0.537\pm 0.177$ $0.500\pm 0.275$ $0.526\pm 0.228$
ViT L1 $0.694\pm 0.136$ $0.616\pm 0.268$ $0.663\pm 0.241$
L2 $0.378\pm 0.129$ $0.516\pm 0.209$ $0.327\pm 0.272$
Fed $0.680\pm 0.138$ $0.648\pm 0.272$ $0.636\pm 0.274$
KD $0.569\pm 0.169$ $0.542\pm 0.248$ $0.566\pm 0.231$
SWIN L1 $0.704\pm 0.138$ $0.647\pm 0.285$ $0.661\pm 0.243$
L2 $0.384\pm 0.199$ $0.519\pm 0.236$ $0.312\pm 0.222$
Fed $0.667\pm 0.155$ $0.652\pm 0.277$ $0.682\pm 0.230$
KD $0.652\pm 0.176$ $0.627\pm 0.273$ $0.670\pm 0.231$